OpenStudy (anonymous):

what is the slope of a perpendicular to x-3y=4 ?

hartnn (hartnn):

can you try to find the slope of x-3y=4 first ?

hartnn (hartnn):

isolate 'y'

hartnn (hartnn):

start by adding 3y on both sides...

OpenStudy (anonymous):

x=4+3y

hartnn (hartnn):

yes so 3y = x-4 divide both sides by 3

OpenStudy (anonymous):

y=x-4/3

OpenStudy (anonymous):

No divide everything by 3

hartnn (hartnn):

y = (x-4)/3 y= x/3 -4/3 compare this with y = mx+c what u get for m ?

OpenStudy (anonymous):

I'm confused a little bit

OpenStudy (anonymous):

So you will get y= 1/3x -4/3 Slope is 1/3 But you need it perpendicular to that slope so flip it over and multiply the slope by -1 So you will get -3

OpenStudy (anonymous):

Perpendicular= flipping it over and multiplying by -1 Parallel=same slope

OpenStudy (anonymous):

thank you both very much :D

OpenStudy (anonymous):

No problem ask me anytime! I'm good with Trigonometry and/or Algebra :)